# Curvature Tuning: Provable Training-free Model Steering From a Single Parameter

### v4

@@ Appendix B Supplementary experimental details
Figure 4: Validation accuracy (%) of CT during the $\beta$ search, averaged over three runs. The accuracy curve varies smoothly and typically peaks in the middle of the $\beta$ range.
B.2 Improving robustness on adversarial and corrupted data with CT (Section 4.2)
B.3 Trainable CT is comparable to LoRA (Section 4.3)
Both Trainable CT and LoRA are trained for 20 epochs using the Adam optimizer. To ensure proper convergence, we use different learning rates: for Trainable CT, a learning rate of $10^{-1}$ is applied to the $(\beta,c)$ parameters and $10^{-3}$ to the linear classifier; for LoRA, a learning rate of $10^{-4}$ is used for both the adapter parameters and the classifier. As before, we apply linear warm-up during the first epoch and decay the learning rate by a factor of 10 after epoch 10.
@@
  • Table 4: mean accuracy over three runs of ImageNet-pretrained ResNet-152 when transferred to 12 downstream datasets, comparing LoRA and Trainable CT.
@@
  • Tables 6 and 7: mean and standard deviation of the learned $\beta$ and $c$ values for Trainable CT across three runs.
@@
  • Figs. 5 and 6: example distributions of $\beta$ and $c$ values in Trainable CT, illustrating commonly and uncommonly observed patterns.
@@
Table 6: Distribution of $\beta$ values in Trainable CT, computed over all $\beta$ parameters across all three runs of ImageNet-pretrained ResNet-18/50/152 and Imagenette-pretrained Swin-T/S when transferred to 12 downstream datasets. The mean and standard deviation of $\beta$ are similar across models (means between 0.69–0.77, stds between 0.31–0.37), suggesting consistent tuning behavior at the model level, while the relatively large standard deviations indicate substantial variation of $\beta$ within each network.
Dataset ResNet-18 ResNet-50 ResNet-152 Swin-T Swin-S
Arabic Characters 0.72 ± 0.34 0.65 ± 0.41 0.68 ± 0.39 0.73 ± 0.35 0.76 ± 0.33
Arabic Digits 0.70 ± 0.43 0.62 ± 0.48 0.62 ± 0.47 0.65 ± 0.42 0.64 ± 0.43
Beans 0.72 ± 0.26 0.76 ± 0.23 0.77 ± 0.19 0.79 ± 0.24 0.83 ± 0.23
CUB-200 0.81 ± 0.17 0.76 ± 0.29 0.79 ± 0.29 0.82 ± 0.27 0.83 ± 0.28
DTD 0.78 ± 0.19 0.77 ± 0.25 0.79 ± 0.24 0.87 ± 0.17 0.88 ± 0.19
FashionMNIST 0.72 ± 0.41 0.65 ± 0.46 0.63 ± 0.46 0.67 ± 0.42 0.66 ± 0.43
FGVC-Aircraft 0.75 ± 0.23 0.70 ± 0.33 0.74 ± 0.32 0.81 ± 0.25 0.82 ± 0.27
Flowers102 0.75 ± 0.16 0.75 ± 0.21 0.79 ± 0.17 0.81 ± 0.22 0.84 ± 0.22
Food101 0.80 ± 0.30 0.71 ± 0.43 0.76 ± 0.40 0.78 ± 0.36 0.74 ± 0.40
DermaMNIST 0.74 ± 0.34 0.70 ± 0.39 0.70 ± 0.37 0.76 ± 0.32 0.77 ± 0.32
OCTMNIST 0.67 ± 0.45 0.62 ± 0.48 0.63 ± 0.47 0.76 ± 0.37 0.64 ± 0.45
PathMNIST 0.69 ± 0.43 0.65 ± 0.47 0.61 ± 0.48 0.78 ± 0.36 0.70 ± 0.43
Average 0.74 ± 0.31 0.69 ± 0.37 0.71 ± 0.35 0.77 ± 0.31 0.76 ± 0.33
@@
Table 7: Distribution of $c$ values in Trainable CT, computed over all $c$ parameters across all three runs of ImageNet-pretrained ResNet-18/50/152 and Imagenette-pretrained Swin-T/S when transferred to 12 downstream datasets. The three ResNet models exhibit similar distributions (means around 0.57–0.59, stds around 0.36–0.38), while the two Swin models also show comparable statistics with higher means (0.67–0.70), and similar stds (0.35-0.37). All models display substantial within-network variation, and the higher average $c$ in Swin models may reflect insufficient pretraining relative to the ResNets.
Dataset ResNet-18 ResNet-50 ResNet-152 Swin-T Swin-S
Arabic Characters 0.63 ± 0.39 0.61 ± 0.39 0.57 ± 0.37 0.66 ± 0.41 0.70 ± 0.38
Arabic Digits 0.59 ± 0.43 0.57 ± 0.42 0.55 ± 0.41 0.63 ± 0.45 0.71 ± 0.43
Beans 0.61 ± 0.29 0.54 ± 0.25 0.53 ± 0.23 0.67 ± 0.26 0.69 ± 0.24
CUB-200 0.60 ± 0.37 0.63 ± 0.37 0.60 ± 0.34 0.70 ± 0.33 0.70 ± 0.33
DTD 0.59 ± 0.31 0.60 ± 0.32 0.57 ± 0.30 0.68 ± 0.25 0.74 ± 0.24
FashionMNIST 0.55 ± 0.44 0.60 ± 0.42 0.56 ± 0.42 0.62 ± 0.46 0.69 ± 0.43
FGVC-Aircraft 0.61 ± 0.36 0.63 ± 0.37 0.58 ± 0.35 0.71 ± 0.33 0.68 ± 0.33
Flowers102 0.58 ± 0.26 0.54 ± 0.26 0.54 ± 0.23 0.65 ± 0.29 0.66 ± 0.25
Food101 0.46 ± 0.47 0.63 ± 0.44 0.60 ± 0.43 0.72 ± 0.42 0.76 ± 0.39
DermaMNIST 0.58 ± 0.38 0.59 ± 0.37 0.57 ± 0.36 0.66 ± 0.36 0.71 ± 0.33
OCTMNIST 0.55 ± 0.45 0.60 ± 0.42 0.57 ± 0.42 0.58 ± 0.47 0.65 ± 0.45
PathMNIST 0.51 ± 0.45 0.58 ± 0.43 0.57 ± 0.42 0.71 ± 0.42 0.76 ± 0.40
Average 0.57 ± 0.38 0.59 ± 0.37 0.57 ± 0.36 0.67 ± 0.37 0.70 ± 0.35
@@
Figure 5: Common distributions of $\beta$ (top) and $c$ (bottom) in Trainable CT across ResNet-18/50/152, averaged over three runs (OCTMNIST shown as a representative dataset). Both $\beta$ and $c$ consistently exhibit sharp U-shaped distributions that appear similar across all models.
@@
Figure 6: Uncommon distributions of $\beta$ (top) and $c$ (bottom) in Trainable CT across ResNet-18/50/152, averaged over three runs (DTD shown as an example dataset). While the overall shape is dataset-specific, the distributions of both $\beta$ and $c$ remain consistent across models.
B.4 CT shows promise on transformers and emerging architectures (Section 4.4)
For the pretraining of Swin-T and Swin-S, we train both models for 200 epochs with a batch size of 128. Training is performed using SGD with a learning rate of 0.1, momentum of 0.9, and weight decay of $5\times 10^{-4}$. We apply linear warm-up during the first epoch and decay the learning rate by a factor of 0.2 at epochs 60, 120, and 160. Training and validation accuracy curves are shown in Fig. 7.
@@
Figure 7: Training and validation accuracy curves for Swin-T and Swin-S during pretraining on Imagenette.

### v5

@@ Appendix B Supplementary experimental details
Figure S2: Validation accuracy (%) of S-CT during the $\beta$ search, averaged over three runs. The accuracy curve varies smoothly and typically peaks in the middle of the $\beta$ range.
Table S4: Mean accuracy (%) ± standard deviation over three runs of ImageNet-pretrained ResNet-18/50/152 and VGG-11 when transferred to 12 downstream datasets. The second row under each method indicates the number of trainable parameters (excluding the linear classifier). By setting $c=0.5$, S-CT performs better than SiLU ($c=1$) and Softplus ($c=0$) on ResNet-18 and ResNet-50, but slightly worse on ResNet-152.
(a) ResNet-18
Dataset Frozen S-CT SiLU Softplus
(0) (1) (1) (1)
Arabic Characters 81.91 ± 0.15 87.65 ± 0.18 81.91 ± 0.19 84.89 ± 0.12
Arabic Digits 97.93 ± 0.08 98.77 ± 0.01 97.95 ± 0.11 98.64 ± 0.06
Beans 87.76 ± 2.05 90.36 ± 1.19 88.80 ± 1.19 88.02 ± 1.19
CUB-200 62.84 ± 0.29 63.18 ± 0.28 62.90 ± 0.33 63.08 ± 0.11
DTD 62.80 ± 0.42 62.66 ± 0.24 62.89 ± 0.33 63.24 ± 0.23
FashionMNIST 88.63 ± 0.13 88.70 ± 0.10 88.63 ± 0.16 88.93 ± 0.09
FGVC-Aircraft 36.80 ± 0.37 38.68 ± 0.05 36.29 ± 0.24 37.44 ± 0.68
Flowers102 80.86 ± 0.29 81.97 ± 0.26 80.86 ± 0.35 83.14 ± 0.03
Food101 61.41 ± 0.07 62.27 ± 0.25 61.37 ± 0.05 62.76 ± 0.06
DermaMNIST 74.83 ± 0.23 75.05 ± 0.60 74.63 ± 0.19 75.13 ± 0.22
OCTMNIST 65.03 ± 0.69 67.27 ± 0.23 65.10 ± 0.82 66.57 ± 0.75
PathMNIST 86.77 ± 0.04 87.51 ± 0.05 86.77 ± 0.05 87.84 ± 0.25
Average 73.96 75.34 74.01 74.97
(b) ResNet-50
Dataset Frozen S-CT SiLU Softplus
(0) (1) (1) (1)
Arabic Characters 80.65 ± 0.07 83.66 ± 0.41 86.46 ± 0.03 81.10 ± 0.21
Arabic Digits 98.33 ± 0.02 98.37 ± 0.06 98.56 ± 0.11 98.41 ± 0.15
Beans 89.58 ± 0.74 91.93 ± 0.90 86.46 ± 1.19 91.15 ± 0.90
CUB-200 65.23 ± 0.43 64.62 ± 0.32 65.12 ± 0.53 65.61 ± 0.50
DTD 67.34 ± 0.16 66.91 ± 0.14 67.27 ± 0.08 67.06 ± 0.62
FashionMNIST 90.05 ± 0.07 90.34 ± 0.23 89.96 ± 0.01 90.35 ± 0.11
FGVC-Aircraft 38.03 ± 0.32 41.16 ± 0.32 38.03 ± 0.39 40.09 ± 0.02
Flowers102 84.00 ± 0.06 83.84 ± 0.13 84.09 ± 0.11 83.94 ± 0.14
Food101 68.06 ± 0.11 68.02 ± 0.11 68.14 ± 0.06 67.90 ± 0.15
DermaMNIST 75.94 ± 0.12 75.89 ± 0.03 75.79 ± 0.12 75.79 ± 0.32
OCTMNIST 67.53 ± 0.21 68.00 ± 0.17 67.33 ± 0.15 67.00 ± 0.20
PathMNIST 90.08 ± 0.22 90.26 ± 0.20 89.95 ± 0.13 89.99 ± 0.14
Average 76.24 76.92 76.43 76.53
(c) ResNet-152
Dataset Frozen S-CT SiLU Softplus
(0) (1) (1) (1)
Arabic Characters 79.86 ± 0.12 79.21 ± 0.55 80.24 ± 0.44 79.83 ± 0.27
Arabic Digits 98.07 ± 0.05 98.15 ± 0.10 98.20 ± 0.18 97.98 ± 0.01
Beans 87.50 ± 1.10 87.50 ± 0.78 88.80 ± 1.97 88.80 ± 2.39
CUB-200 67.68 ± 0.54 68.15 ± 0.62 67.68 ± 0.66 69.80 ± 0.32
DTD 66.95 ± 0.03 66.97 ± 0.05 66.86 ± 0.23 66.95 ± 0.03
FashionMNIST 90.37 ± 0.11 90.44 ± 0.16 90.28 ± 0.26 90.60 ± 0.03
FGVC-Aircraft 38.74 ± 0.16 38.51 ± 0.14 38.74 ± 0.20 39.20 ± 0.38
Flowers102 82.98 ± 0.16 83.28 ± 0.25 82.97 ± 0.19 83.54 ± 0.13
Food101 71.11 ± 0.09 71.13 ± 0.08 71.11 ± 0.11 71.19 ± 0.08
DermaMNIST 75.68 ± 0.47 76.23 ± 0.14 74.76 ± 0.69 76.16 ± 0.10
OCTMNIST 69.27 ± 0.98 69.10 ± 1.47 69.47 ± 1.29 69.90 ± 0.50
PathMNIST 89.91 ± 0.12 89.82 ± 0.09 89.91 ± 0.15 90.75 ± 0.06
Average 76.51 76.54 76.59 77.06
B.2 T-CT is comparable to LoRA with fewer parameters (Section˜4.2)
Both T-CT and LoRA are trained for 20 epochs using the Adam optimizer. To ensure proper convergence, we use different learning rates: for T-CT, a learning rate of $10^{-1}$ is applied to the $(\beta,c)$ parameters and $10^{-3}$ to the linear classifier; for LoRA, a learning rate of $10^{-4}$ is used for both the adapter parameters and the classifier. As before, we apply linear warm-up during the first epoch and decay the learning rate by a factor of 10 after epoch 10.
@@
  • Table˜S1: mean accuracy (± std) over three runs of ImageNet-pretrained ResNet-18/50/152 and VGG-11 when transferred to 12 downstream datasets, comparing LoRA and T-CT.
  • Table˜S5: additional experiments on LoRA with rank $r\in\{1,2,4\}$ and scaling factors $\alpha\in\{r,2r,4r\}$.
@@
  • Tables˜S6 and S7: mean (± std) of the learned $\beta$ and $c$ values of T-CT across three runs.
@@
  • Figs.˜S3 and S4: example distributions of $\beta$ and $c$ values in T-CT, illustrating commonly and uncommonly observed patterns.
Table S5: Mean accuracy (%) ± standard deviation over three runs of ImageNet-pretrained ResNet-18/50/152 when transferred to 12 downstream datasets. The second row under each method indicates the number of trainable parameters (excluding the linear classifier). For each dataset and LoRA rank $r$, we report the best test accuracy achieved among the candidate scaling factors $\alpha\in\{r,2r,4r\}$. T-CT can still outperform LoRA under this more challenging setting.
(a) ResNet-18
Dataset T-CT LoRA ($r=1$) LoRA ($r=2$) LoRA ($r=4$)
(3968) (35923) (71846) (143692)
Arabic Characters 93.76 ± 0.22 94.23 ± 0.13 95.30 ± 0.18 96.26 ± 0.12
Arabic Digits 99.03 ± 0.01 99.12 ± 0.02 99.21 ± 0.05 99.23 ± 0.03
Beans 94.01 ± 0.37 94.01 ± 0.74 95.83 ± 0.37 95.83 ± 1.33
CUB-200 64.30 ± 0.16 54.83 ± 0.37 56.11 ± 0.26 57.47 ± 0.51
DTD 63.62 ± 0.67 54.36 ± 0.31 56.17 ± 0.22 57.93 ± 0.43
FashionMNIST 91.07 ± 0.16 92.03 ± 0.11 92.83 ± 0.04 93.50 ± 0.05
FGVC-Aircraft 46.44 ± 0.49 29.50 ± 0.92 32.94 ± 0.40 39.13 ± 0.32
Flowers102 86.55 ± 0.21 67.53 ± 0.76 69.68 ± 0.89 73.30 ± 0.46
Food101 66.04 ± 0.17 64.40 ± 0.08 65.96 ± 0.32 66.97 ± 0.35
DermaMNIST 77.66 ± 0.29 75.54 ± 0.20 76.79 ± 0.77 76.72 ± 0.51
OCTMNIST 69.53 ± 1.11 74.83 ± 0.17 76.37 ± 0.66 76.47 ± 0.26
PathMNIST 87.17 ± 0.66 87.78 ± 0.18 88.08 ± 1.15 88.85 ± 0.48
Average 78.26 74.01 75.44 76.80
(b) ResNet-50
Dataset T-CT LoRA ($r=1$) LoRA ($r=2$) LoRA ($r=4$)
(45440) (79443) (158886) (317772)
Arabic Characters 95.67 ± 0.03 94.38 ± 0.22 95.67 ± 0.34 96.30 ± 0.04
Arabic Digits 99.16 ± 0.03 99.09 ± 0.01 99.22 ± 0.09 99.22 ± 0.02
Beans 95.57 ± 0.74 96.35 ± 0.37 96.09 ± 1.69 95.31 ± 0.64
CUB-200 71.03 ± 0.64 66.17 ± 0.51 67.91 ± 0.53 68.93 ± 0.23
DTD 65.07 ± 0.37 64.79 ± 0.30 64.91 ± 0.49 67.07 ± 0.31
FashionMNIST 92.78 ± 0.06 92.19 ± 0.17 92.90 ± 0.14 93.59 ± 0.13
FGVC-Aircraft 55.70 ± 0.76 42.12 ± 0.17 47.46 ± 0.19 52.64 ± 0.47
Flowers102 87.62 ± 0.28 82.58 ± 0.47 83.39 ± 0.35 84.63 ± 0.29
Food101 73.60 ± 0.13 71.42 ± 0.14 73.01 ± 0.24 74.89 ± 0.05
DermaMNIST 78.02 ± 0.50 76.21 ± 0.27 77.26 ± 0.36 77.39 ± 0.47
OCTMNIST 74.13 ± 1.65 76.23 ± 0.09 76.07 ± 1.19 77.83 ± 0.82
PathMNIST 87.33 ± 0.74 87.29 ± 0.67 86.04 ± 0.23 87.44 ± 0.16
Average 81.31 79.07 79.99 81.27
(c) ResNet-152
Dataset T-CT LoRA ($r=1$) LoRA ($r=2$) LoRA ($r=4$)
(143744) (243283) (486566) (973132)
Arabic Characters 96.47 ± 0.39 96.00 ± 0.16 96.43 ± 0.02 96.87 ± 0.11
Arabic Digits 99.10 ± 0.05 99.16 ± 0.03 99.21 ± 0.03 99.25 ± 0.01
Beans 96.35 ± 1.33 94.53 ± 1.10 97.92 ± 0.37 97.14 ± 0.74
CUB-200 73.04 ± 0.19 70.75 ± 0.59 70.94 ± 0.15 71.72 ± 0.43
DTD 63.39 ± 0.34 66.63 ± 0.07 67.66 ± 0.50 68.28 ± 0.51
FashionMNIST 93.39 ± 0.12 92.77 ± 0.04 93.49 ± 0.04 93.98 ± 0.14
FGVC-Aircraft 58.16 ± 0.31 49.06 ± 0.26 55.82 ± 1.04 59.98 ± 0.26
Flowers102 83.43 ± 1.01 84.51 ± 0.58 84.83 ± 0.17 86.24 ± 0.04
Food101 76.08 ± 0.15 74.66 ± 0.08 76.00 ± 0.16 76.86 ± 0.10
DermaMNIST 77.94 ± 0.60 77.02 ± 0.69 77.31 ± 0.75 77.46 ± 0.16
OCTMNIST 75.17 ± 2.10 77.90 ± 0.36 78.23 ± 1.32 78.63 ± 0.21
PathMNIST 83.60 ± 0.42 86.75 ± 0.86 88.33 ± 0.33 86.81 ± 1.95
Average 81.34 80.81 82.18 82.77
@@
Table S6: Distribution of $\beta$ values in T-CT, computed over all $\beta$ parameters across all three runs of ImageNet-pretrained ResNet-18/50/152 and VGG-11 when transferred to 12 downstream datasets. The mean and standard deviation of $\beta$ are similar across models (means between 0.69–0.79, stds between 0.31–0.37), suggesting consistent tuning behavior at the model level, while the relatively large standard deviations indicate substantial variation of $\beta$ within each network.
Dataset ResNet-18 ResNet-50 ResNet-152 VGG-11
Arabic Characters 0.72 ± 0.34 0.65 ± 0.41 0.68 ± 0.39 0.73 ± 0.39
Arabic Digits 0.70 ± 0.43 0.62 ± 0.48 0.62 ± 0.47 0.73 ± 0.43
Beans 0.72 ± 0.26 0.76 ± 0.23 0.77 ± 0.19 0.79 ± 0.20
CUB-200 0.81 ± 0.17 0.76 ± 0.29 0.79 ± 0.29 0.75 ± 0.31
DTD 0.78 ± 0.19 0.77 ± 0.25 0.79 ± 0.24 0.80 ± 0.24
FashionMNIST 0.72 ± 0.41 0.65 ± 0.46 0.63 ± 0.46 0.81 ± 0.37
FGVC-Aircraft 0.75 ± 0.23 0.70 ± 0.33 0.74 ± 0.32 0.74 ± 0.31
Flowers102 0.75 ± 0.16 0.75 ± 0.21 0.79 ± 0.17 0.74 ± 0.22
Food101 0.80 ± 0.30 0.71 ± 0.43 0.76 ± 0.40 0.88 ± 0.27
DermaMNIST 0.74 ± 0.34 0.70 ± 0.39 0.70 ± 0.37 0.81 ± 0.33
OCTMNIST 0.67 ± 0.45 0.62 ± 0.48 0.63 ± 0.47 0.83 ± 0.35
PathMNIST 0.69 ± 0.43 0.65 ± 0.47 0.61 ± 0.48 0.82 ± 0.37
Average 0.74 ± 0.31 0.69 ± 0.37 0.71 ± 0.35 0.79 ± 0.32
@@
Table S7: Distribution of $c$ values in T-CT, computed over all $c$ parameters across all three runs of ImageNet-pretrained ResNet-18/50/152 and VGG-11 when transferred to 12 downstream datasets. The mean and standard deviation of $c$ are similar across models (means between 0.57–0.61, stds between 0.32–0.38), suggesting consistent tuning behavior at the model level, while the relatively large standard deviations indicate substantial variation of $c$ within each network.
Dataset ResNet-18 ResNet-50 ResNet-152 VGG-11
Arabic Characters 0.63 ± 0.39 0.61 ± 0.39 0.57 ± 0.37 0.65 ± 0.33
Arabic Digits 0.59 ± 0.43 0.57 ± 0.42 0.55 ± 0.41 0.62 ± 0.38
Beans 0.61 ± 0.29 0.54 ± 0.25 0.53 ± 0.23 0.55 ± 0.21
CUB-200 0.60 ± 0.37 0.63 ± 0.37 0.60 ± 0.34 0.67 ± 0.31
DTD 0.59 ± 0.31 0.60 ± 0.32 0.57 ± 0.30 0.59 ± 0.27
FashionMNIST 0.55 ± 0.44 0.60 ± 0.42 0.56 ± 0.42 0.61 ± 0.38
FGVC-Aircraft 0.61 ± 0.36 0.63 ± 0.37 0.58 ± 0.35 0.66 ± 0.31
Flowers102 0.58 ± 0.26 0.54 ± 0.26 0.54 ± 0.23 0.60 ± 0.22
Food101 0.46 ± 0.47 0.63 ± 0.44 0.60 ± 0.43 0.58 ± 0.40
DermaMNIST 0.58 ± 0.38 0.59 ± 0.37 0.57 ± 0.36 0.56 ± 0.30
OCTMNIST 0.55 ± 0.45 0.60 ± 0.42 0.57 ± 0.42 0.62 ± 0.39
PathMNIST 0.51 ± 0.45 0.58 ± 0.43 0.57 ± 0.42 0.58 ± 0.40
Average 0.57 ± 0.38 0.59 ± 0.37 0.57 ± 0.36 0.61 ± 0.32
@@
Figure S3: Common distributions of $\beta$ (top) and $c$ (bottom) in T-CT across ResNet-18/50/152, averaged over three runs (OCTMNIST shown as a representative dataset). Both $\beta$ and $c$ consistently exhibit sharp U-shaped distributions that appear similar across all models.
@@
Figure S4: Uncommon distributions of $\beta$ (top) and $c$ (bottom) in T-CT across ResNet-18/50/152, averaged over three runs (DTD shown as an example dataset). While the overall shape is dataset-specific, the distributions of both $\beta$ and $c$ remain consistent across models.
B.3 Improving model robustness through CT’s implicit bias (Section˜4.3)
Additional results are provided as follows:
  • Table˜S8: mean robust accuracy (± std) over three runs of ImageNet-pretrained ResNet-18/50/152 under $\ell_{\infty}$$\ell_{2}$ attacks and corruptions on CIFAR-10/100 and ImageNet.
  • Table˜S9: mean robust accuracy (± std) over three runs of ImageNet-pretrained ResNet-18/50/152 transferred to CIFAR-10/100 under $\ell_{\infty}$, $\ell_{2}$ attacks, and corruptions.
Table S8: Mean robust accuracy (%) ± standard deviation over three runs of ImageNet-pretrained ResNet-18/50/152 under $\ell_{2}$$\ell_{\infty}$ attacks and corruptions on CIFAR-10/100 and ImageNet. S-CT yields substantial improvements under $\ell_{\infty}$ attacks, with the selected $\beta$ values close to 1.
Robustness Model Dataset Frozen S-CT $\beta$
$\ell_{2}$ ResNet-18 CIFAR-10 53.67 ± 0.32 53.67 ± 0.32 1.00 ± 0.00
CIFAR-100 24.30 ± 0.10 25.50 ± 0.00 0.92 ± 0.00
ImageNet 23.37 ± 0.06 23.37 ± 0.06 1.00 ± 0.00
Average 33.78 34.18 0.97
ResNet-50 CIFAR-10 55.10 ± 0.10 56.53 ± 0.21 0.97 ± 0.00
CIFAR-100 23.83 ± 0.06 25.80 ± 0.20 0.96 ± 0.00
ImageNet 31.90 ± 0.00 31.90 ± 0.00 1.00 ± 0.00
Average 36.94 38.08 0.98
ResNet-152 CIFAR-10 56.27 ± 0.23 56.27 ± 0.23 1.00 ± 0.00
CIFAR-100 27.90 ± 0.10 28.23 ± 0.12 0.98 ± 0.00
ImageNet 42.50 ± 0.00 42.50 ± 0.00 1.00 ± 0.00
Average 42.22 42.33 0.99
$\ell_{\infty}$ ResNet-18 CIFAR-10 11.17 ± 0.06 14.93 ± 0.06 0.90 ± 0.00
CIFAR-100 4.47 ± 0.06 6.90 ± 0.00 0.92 ± 0.00
ImageNet 0.00 ± 0.00 7.00 ± 0.10 0.89 ± 0.00
Average 5.21 9.61 0.90
ResNet-50 CIFAR-10 10.10 ± 0.17 14.83 ± 0.06 0.95 ± 0.00
CIFAR-100 4.43 ± 0.06 7.90 ± 0.00 0.93 ± 0.00
ImageNet 0.30 ± 0.00 9.30 ± 0.17 0.93 ± 0.00
Average 4.94 9.76 0.94
ResNet-152 CIFAR-10 11.47 ± 0.06 15.00 ± 0.20 0.99 ± 0.00
CIFAR-100 5.40 ± 0.00 7.70 ± 0.17 0.99 ± 0.00
ImageNet 0.30 ± 0.00 13.53 ± 0.06 0.97 ± 0.01
Average 5.72 12.08 0.98
Corruptions ResNet-18 CIFAR-10 77.73 ± 0.00 77.73 ± 0.00 1.00 ± 0.00
CIFAR-100 51.81 ± 0.00 51.95 ± 0.00 0.94 ± 0.00
ImageNet 33.11 ± 0.00 33.32 ± 0.00 0.92 ± 0.00
Average 54.22 54.33 0.95
ResNet-50 CIFAR-10 77.26 ± 0.00 77.26 ± 0.00 1.00 ± 0.00
CIFAR-100 53.91 ± 0.00 53.93 ± 0.00 0.98 ± 0.00
ImageNet 39.64 ± 0.00 39.64 ± 0.00 1.00 ± 0.00
Average 56.94 56.94 0.99
ResNet-152 CIFAR-10 78.82 ± 0.00 78.83 ± 0.00 0.99 ± 0.00
CIFAR-100 56.12 ± 0.00 56.12 ± 0.00 1.00 ± 0.00
ImageNet 45.47 ± 0.00 45.47 ± 0.00 0.99 ± 0.00
Average 60.14 60.14 0.99
Table S9: Mean robust accuracy (%) ± standard deviation over three runs of ImageNet-pretrained ResNet-18/50/152 transferred to CIFAR-10/100 under $\ell_{2}$, $\ell_{\infty}$ attacks, and corruptions. T-CT improves $\ell_{\infty}$ robustness significantly compared to linear probing and LoRA.
Robustness Model Dataset Frozen LoRA T-CT
$\ell_{2}$ ResNet18 CIFAR10 8.47 ± 0.26 5.93 ± 1.65 8.93 ± 0.37
CIFAR100 1.57 ± 0.21 0.77 ± 0.33 1.10 ± 0.45
Average 5.02 3.35 5.01
ResNet50 CIFAR10 6.23 ± 0.34 4.57 ± 1.32 6.83 ± 1.48
CIFAR100 0.70 ± 0.08 0.37 ± 0.26 0.47 ± 0.31
Average 3.47 2.47 3.65
ResNet152 CIFAR10 8.03 ± 0.52 4.63 ± 2.01 8.00 ± 1.22
CIFAR100 0.90 ± 0.08 0.47 ± 0.26 0.50 ± 0.08
Average 4.46 2.55 4.25
$\ell_{\infty}$ ResNet18 CIFAR10 0.30 ± 0.00 0.70 ± 0.71 1.57 ± 0.74
CIFAR100 0.03 ± 0.05 0.07 ± 0.05 0.17 ± 0.12
Average 0.16 0.38 0.87
ResNet50 CIFAR10 0.20 ± 0.08 0.33 ± 0.29 2.43 ± 1.54
CIFAR100 0.00 ± 0.00 0.03 ± 0.05 0.07 ± 0.09
Average 0.10 0.18 1.25
ResNet152 CIFAR10 0.43 ± 0.09 0.20 ± 0.14 5.10 ± 2.97
CIFAR100 0.17 ± 0.05 0.00 ± 0.00 0.00 ± 0.00
Average 0.30 0.10 2.55
Corruptions ResNet18 CIFAR10 21.34 ± 0.29 13.59 ± 0.30 16.83 ± 2.36
CIFAR100 5.10 ± 0.15 2.96 ± 1.05 4.62 ± 0.68
Average 13.22 8.28 10.72
ResNet50 CIFAR10 16.23 ± 0.21 11.69 ± 0.90 12.68 ± 2.06
CIFAR100 3.47 ± 0.09 2.04 ± 0.36 1.61 ± 0.13
Average 9.85 6.86 7.14
ResNet152 CIFAR10 13.82 ± 0.49 11.33 ± 1.22 9.83 ± 2.07
CIFAR100 2.07 ± 0.12 2.13 ± 0.22 1.72 ± 0.51
Average 7.94 6.73 5.78
B.4 CT shows promise on transformers (Section˜4.4)
In this experiment, LoRA is applied to all QKV projection layers. For all three methods—linear probing, LoRA, and T-CT—we perform a grid search over the learning rate and select the model achieving the best validation accuracy for testing. The learning rate is selected from $10^{-2},10^{-3},10^{-4}$} for linear probing, from $10^{-3},10^{-4},10^{-5}$} for LoRA, and for T-CT, we fix the learning rate of the linear classifier to $10^{-3}$ while searching over $10^{-1},10^{-2},10^{-3},10^{-4}$} for the $(\beta,c)$ parameters. All other training configurations follow those described in Sections˜B.1 and B.2.
Additional results are provided as follows:
  • Table˜S10: mean accuracy (± std) over three runs of ImageNet-pretrained Swin-T/S when transferred to 12 downstream datasets.
Table S10: Mean accuracy (%) ± standard deviation over three runs of ImageNet-pretrained Swin-T/S when transferred to 12 downstream datasets. The second row under each method indicates the number of trainable parameters (excluding the linear classifier). T-CT improves over linear probing but underperforms LoRA.
@@
Dataset Frozen LoRA T-CT
(0) (74832) (532)
Arabic Characters 83.48 ± 0.15 93.24 ± 0.13 85.02 ± 0.30
Arabic Digits 98.14 ± 0.07 99.19 ± 0.01 98.47 ± 0.04
Beans 88.28 ± 1.10 94.01 ± 0.37 89.06 ± 1.10
CUB-200 73.42 ± 0.17 78.73 ± 0.28 74.33 ± 0.14
DTD 70.66 ± 0.13 70.99 ± 0.61 71.45 ± 0.31
FashionMNIST 89.89 ± 0.04 93.15 ± 0.13 90.23 ± 0.03
FGVC-Aircraft 48.06 ± 0.32 48.29 ± 0.46 47.58 ± 0.99
Flowers102 86.66 ± 0.17 90.22 ± 0.34 85.35 ± 0.20
Food101 77.05 ± 0.03 83.69 ± 0.11 78.90 ± 0.11
DermaMNIST 75.83 ± 0.27 76.71 ± 0.43 75.86 ± 0.11
OCTMNIST 69.97 ± 0.62 76.30 ± 1.66 67.97 ± 1.01
PathMNIST 89.14 ± 0.23 92.26 ± 0.12 91.73 ± 0.19
Average 77.69 82.23 78.53
@@
Dataset Frozen LoRA T-CT
(0) (148560) (868)
Arabic Characters 83.83 ± 0.05 94.38 ± 0.34 86.65 ± 0.50
Arabic Digits 98.28 ± 0.03 99.19 ± 0.05 98.39 ± 0.04
Beans 90.89 ± 1.95 95.05 ± 1.47 91.41 ± 0.64
CUB-200 72.66 ± 0.56 79.45 ± 0.52 73.40 ± 0.10
DTD 69.77 ± 0.44 71.56 ± 0.66 72.43 ± 0.13
FashionMNIST 89.75 ± 0.03 93.52 ± 0.05 89.85 ± 0.08
FGVC-Aircraft 44.36 ± 0.21 51.94 ± 0.60 45.72 ± 0.27
Flowers102 83.24 ± 0.05 87.67 ± 3.41 85.08 ± 0.25
Food101 77.59 ± 0.06 85.17 ± 0.23 79.45 ± 0.14
DermaMNIST 76.64 ± 0.22 78.15 ± 0.67 77.14 ± 0.02
OCTMNIST 66.90 ± 0.29 76.97 ± 0.45 69.07 ± 0.60
PathMNIST 89.74 ± 0.38 92.79 ± 0.33 92.13 ± 0.15
Average 78.06 82.90 79.24